Help Needed;
Multiply both sides of each equation by a common denominator to eliminate the fractions. Then solve the system.
3x - ((5y)/(7)) = 10
((2x)/(3) - ((9y)/(7)) = (19/5)

Answer by profemmanuel2q@yahoo.com(15)

3x - ((5y)/(7)) = 10 - Eqn. 1
((2x)/(3) - ((9y)/(7)) = (19/5) - Eqn 2
Taking Equation 1, Multiply both sides by the L.C.M (Leaset Common Multiple, that is 7)so that we will be able to get rid of the denominator.
7 X 3x - 7 X 5y/7 = 7 X 10
21x - 5y = 70 - Eqn. 1
Taking Equation 2,
Multiply both sides of the L.C.M, that is 105, so we have
105 X 2x/3 - 105 X 9y/7 = 105 X 19/5, Canceling the
35 X 2x - 15 X 9y = 21 X 19
70x - 135y = 399 - Eqn. 2
Solving Equation 1 and Equation 2 simultaneously:
21x - 5y = 70 - Eqn 1
70x -135y = 399 - Eqn 2
Before you can solve this: You have to choose the method type of solving simultaneous Equation. Whether using the Elimination method, Substitution method or the Graphical method.
When u solve this you are suppose to get x = -3.64 and y = -17.472